OpenStudy (anonymous):

What is the value of the coefficient "a" when the quadratic equation y = (3x − 4)(2x − 1) is written in standard form?

OpenStudy (tommo_lcfc):

y = ax^2 + bx + c = 6x^2 - 11x + 4. So a = 6